OpenStudy (anonymous):

50g^2+40g+8 factor out

OpenStudy (anonymous):

50 *8=400 20*20=400 50g^2+20g+20g+8=0 20g(2.5g+1)+8 (2.5g+1)=0 20g+8=0 and 2.5g+1=0 20g=-8 g=-8/20 2.5g=-1 g=-0.4